### v1

# Exploring Equity of Climate Policies using Multi-Agent Multi-Objective Reinforcement Learning

Abstract
Addressing climate change requires coordinated policy efforts of nations worldwide. These efforts are informed by scientific reports, which rely in part on Integrated Assessment Models (IAMs), prominent tools used to assess the economic impacts of climate policies. However, traditional IAMs optimize policies based on a single objective, limiting their ability to capture the trade-offs among economic growth, temperature goals, and climate justice. As a result, policy recommendations have been criticized for perpetuating inequalities, fueling disagreements during policy negotiations. We introduce Justice, the first framework integrating IAM with Multi-Objective Multi-Agent Reinforcement Learning (MOMARL). By incorporating multiple objectives, Justice generates policy recommendations that shed light on equity while balancing climate and economic goals. Further, using multiple agents can provide a realistic representation of the interactions among the diverse policy actors. We identify equitable Pareto-optimal policies using our framework, which facilitates deliberative decision-making by presenting policymakers with the inherent trade-offs in climate and economic policy.
1 Introduction
Climate change poses a significant global threat, disproportionately affecting marginalized communities (Faus Onbargi, 2022; Rising et al., 2022). The extent of this threat remains uncertain due to the complex interplay between climate and socioeconomic systems (Burke et al., 2018; van der Wijst et al., 2021). This uncertainty underscores a central challenge of climate justice: the fair allocation of burdens and benefits through the implementation of climate policies (Pozo et al., 2020). The complexity in assessing the implications of policy measures leads to contentious international negotiations, as seen in the Conference of the Parties (COP). Asymmetric impacts, divergent responsibilities, and differing national priorities often result in impasses and heated disagreements among policymakers (Wei et al., 2013; Lamb et al., 2020). Climate change is thus characterized as a wicked problem in the policy domain (Lönngren and Svanström, 2016; Verweij, 2023), necessitating interventions that seek to balance the diverse ethical preferences of stakeholders.
The Intergovernmental Panel on Climate Change (IPCC) is a key player in the global climate change discourse, and its assessment reports serve as the primary source of scientific information for international climate negotiations, such as the COP (Asayama, 2024). These technical reports, crafted by experts from various disciplines, including climate science and economics, provide evidence-based guidance to support effective climate action through the use of Integrated Assessment Models (IAMs) (Cointe et al., 2019). IAMs integrate socioeconomic, climate and technological processes into a single framework to assess global mitigation pathways and simulate future socio-economic and climate scenarios.
Although IAMs are influential tools, they are criticized for suggesting inequitable mitigation policies that disadvantage developing countries (Rivadeneira and Carton, 2022; Gambhir et al., 2022). These models often simplify the complexity of climate change policies to a single objective, which is ethically problematic, as it favours dominant stakeholders and neglects non-economic metrics, such as temperature, biodiversity, and human mortality (Bromley and Beattie, 1973; Kasprzyk et al., 2016). Recent studies seek to address these shortcomings by enhancing IAMs with multiple objectives (Marangoni et al., 2021; Ferrari et al., 2022), or by incorporating multi-agent approaches, such as multi-regional studies (Zhang et al., 2022). However, the multi-objective studies treat the world as a single agent, which overlooks the interests of various stakeholders. Conversely, the multi-agent approach focuses on a single objective, neglecting the diverse preferences of agents during negotiations. Currently, no single modeling framework simultaneously addresses both the multiplicity of objectives and agents, thus overlooking the complexities of the real-world. The exclusion of either multiple objectives or multiple agents limits the applicability of IAMs. Our contribution fills this gap by developing Justice, a Multi-Objective Multi-Agent Reinforcement Learning (MOMARL) IAM framework that emulates real-world negotiations and discovers equitable policy options.  [^1] Code: https://anonymous.4open.science/r/justice-rl-7CC5
MOMARL is a powerful framework for complex decision-making that requires balancing conflicting objectives and coordinating independent decisions over time. In particular, it is a crucial tool for addressing problems that involve sequential decisions (Radulescu, 2024). By extending reinforcement learning (RL) to accommodate multiple agents and multiple objectives, MOMARL enables trade-off management through vectorial rewards, where each component reflects performance on a specific objective. Despite the prevalence of societal problems involving multiple stakeholders with diverse preferences, the field of MOMARL remains underexplored. Simplifications such as hard-coding trade-offs or centralizing decisions limit practical applicability. Importantly, existing works are limited to theoretical examples or simple (grid-like) environments (Felten et al., 2024b; Rădulescu et al., 2020).
Our contribution is twofold. (1) For the IAM community, Justice introduces a multi-objective, multi-agent framework of a climate-economy model, providing a tool that can inform IPCC’s synthesis reports. Such a framework can enable the design of equitable policies that account for regional disparities while balancing multiple aspects of climate change, including economic and environmental outcomes. (2) For the MOMARL community, Justice offers a well-designed, open-source, implementation using the MOMALand API (Felten et al., 2024b), enabling seamless integration with any RL algorithm. This provides a real-world testbed for evaluating and benchmarking future algorithms, supporting the development of robust and societal applications of reinforcement learning methods.
2 Background
Our interdisciplinary work is based on two foundational areas: climate modelling and reinforcement learning.
2.1 Integrated Assessment Models
IAMs offer a holistic approach to inform policy decisions on climate change (Gambhir et al., 2019). These models integrate socioeconomic, technological, and biogeochemical variables to represent interactions between simplified social and climate components (Rivadeneira and Carton, 2022). IAMs primarily inform climate mitigation policies and are frequently used for policy evaluation and optimization (Mastrandrea, 2009). Originating from William Nordhaus’s DICE model [1992], which combined economic and climate models for global cost-benefit analyses of climate policies, IAMs analyze climate policy’s impacts on both the economy and environment. DICE is a foundational framework, earning Nordhaus a Nobel Prize and being utilized by the US government to calculate the social cost of carbon, which measures the societal benefit of reducing CO2 emissions (Grubb et al., 2021). DICE has inspired numerous IAMs and remains a valuable tool for assessing climate policies. A popular regional variation of DICE, known as RICE (Nordhaus and Yang, 1996), encompasses 12 regions, and its recent and enhanced version called RICE50+ (Gazzotti et al., 2021) expands to 57 regions for greater regional resolution.
IAMs can be broadly categorized into two types: (1) highly aggregated Cost-Benefit models and (2) detailed process-based models (Van Beek et al., 2020). Cost-Benefit IAMs (CB-IAMs) are optimization models that identify near-term emission reduction pathways to maximize long-term benefits by considering high-level economic and climate interactions. DICE/RICE IAMs fall under this CB-optimization category. In contrast, Process-Based IAMs (PB-IAMs) are simulation models that offer detailed economic representations across multiple sectors to analyze the impacts of specific policies on economic, social, and environmental factors, with an emphasis on sector-specific environmental impacts, such as those in energy systems (Nikas et al., 2019). CB-IAMs are employed for global mitigation pathways, macroeconomic assessments of mitigation strategies, and strategic interactions reported by the IPCC’s Working Group III (WGIII). Notably, the macroeconomic components of CB-IAMs form the foundation of many process-based models. The global mitigation pathways found by CB-IAMs can also be used to inform the simulation of PB-IAMs. The DICE/RICE family, despite its simplicity, has significantly influenced IPCC WGIII mitigation assessments in various reports, including the recent sixth assessment report, particularly in Chapters 3 and 14, which focus on Mitigation Pathways and International Cooperation, respectively (Shukla et al., 2022).
2.2 Multi-Objective Multi-Agent RL
We formalize a MOMARL problem as a multi-objective multi-agent Markov decision process (MOMAMDP) with team reward (Rădulescu et al., 2020), defined as a tuple $\bigl{(}\mathcal{S},\mathcal{A},F,\mathbf{R}\bigr{)}$, comprising $N\geq 2$ agents and $d\geq 2$ objectives:
  • $S$ is the state space.
  • $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}$ is the set of joint actions, where $\mathcal{A}_{n}$ denotes the action set of agent $n$.
  • $F:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the probabilistic transition function, mapping each state–joint action pair to a distribution over next states.
  • $\mathbf{R}=\mathbf{R}_{1}\times\cdots\times\mathbf{R}_{n}$ represents the reward functions, where $\mathbf{R}_{n}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ is the vector-valued reward function for agent $n$ across $d$ objectives.
Agents optimize policies $\pi_{n}$ to maximize expected discounted returns:
$\mathbf{v}^{\pi}_{n}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbf{R}_{n}(s_{t},\mathbf{a}_{t},s_{t+1})\mid\boldsymbol{\pi}\right]$
where $\boldsymbol{\pi}=({\pi}_{1},\dots,{\pi}_{n})$ is the joint vector policy of the agents acting in the environment, $\gamma$ is the discount factor, and $\mathbf{R}_{n}(s_{t},\mathbf{a}_{t},s_{t+1})$ is the vectorial reward obtained by agent $n$ at timestep $t$ for the joint action $\mathbf{a}_{t}\in\mathcal{A}$ at state $s_{t}\in\mathcal{S}$.
Solution Set Concept
The value function $\mathbf{v}^{\boldsymbol{\pi}}_{n}\in\mathbb{R}^{d}$ offers a partial ordering over policies, as it is a vector. Identifying the optimal policy requires a utility function $u_{n}:\mathbb{R}^{d}\to\mathbb{R}$ to capture agents’ preferences by mapping vectors to scalars.
The Pareto set, assuming the utility function is monotonically increasing, uses Pareto dominance ($\succ_{P}$), where a vector dominates another if it is at least equal in all objectives and strictly better in one. When agents share a team reward ($\mathbf{v}_{1}^{\boldsymbol{\pi}}=\mathbf{v}_{2}^{\boldsymbol{\pi}}=\cdots=\mathbf{v}_{n}^{\boldsymbol{\pi}}=\mathbf{v}^{\boldsymbol{\pi}}$), Pareto dominance applies directly. For a set of policies $\boldsymbol{\Pi}$, the Pareto set $P(\boldsymbol{\Pi})$ includes all undominated policies:
$P(\boldsymbol{\Pi})=\{\boldsymbol{\pi}\in\boldsymbol{\Pi}\mid\nexists\boldsymbol{\pi}^{\prime}\in\boldsymbol{\Pi}:\mathbf{v}^{\boldsymbol{\pi}^{\prime}}\succ_{P}\mathbf{v}^{\boldsymbol{\pi}}\}.$
The Pareto front $F(P)$ contains value vectors for Pareto-optimal policies in $P(\boldsymbol{\Pi})$. When the utility function is a positively weighted linear sum, the solution set forms the convex hull of $\mathbf{v}^{\boldsymbol{\pi}}$ (Felten et al., 2024a).
3 The Justice Model
The Justice IAM is a modular and efficient framework created to assess various modeling assumptions related to economic growth, damage functions, abatement costs, and social welfare. It allows experimentation with different uncertainties without incurring high computational costs. We provide an outline of the model and its MOMARL formulation.
3.1 High-Level Outline
Justice incorporates the economy, damage, and abatement modules from the RICE50+ model (Gazzotti et al., 2021) and integrates them with the FAIR climate model (Smith et al., 2023), facilitating probabilistic assessments of climate policies across various climate sensitivity scenarios. FAIR is a simplified emulator of complex climate models grounded in climate physics that accurately reproduces historical climate data and is utilized by the IPCC(Leach et al., 2021; Shukla et al., 2022). Unlike other IAMs, Justice identifies Pareto-optimal policies in a multi-objective framework, assessing their robustness across various socioeconomic and climate uncertainties and considering the distributional impacts of the Pareto-optimal policies.
Justice operates on a yearly resolution (instead of a 5-year interval used in RICE50+) and covers 57 independent regions (Table LABEL:tab:justice_macro_regions_grouped in Appendix A [^2] Online Appendix: https://anonymous.4open.science/r/justice-rl-7CC5/Submission_Appendix.pdf). Justice utilizes the Shared Socioeconomic Pathway (SSP) dataset (Riahi et al., 2017) for exogenous regional economic growth, carbon intensity, and population projections, and the Representative Concentration Pathway (RCP) dataset (Meinshausen et al., 2011) informs emission trajectories for other greenhouse gases and aerosols. The SSPs and RCPs provide a consistent framework for integrating socioeconomic and climate research, representing various future scenarios through different socioeconomic narratives and climate forcings. The exogenous SSP data is denoted with a superscript $(^{*})$ in the equations.
Figure 1 provides an overview of the Justice IAM. The economy sub-model employs the neoclassical Cobb-Douglas production function (Gazzotti et al., 2021), as in the DICE IAM, to compute economic growth using exogenous labour and total factor productivity data along with capital stock. The resulting economic output feeds the Emissions sub-model, which calculates CO2 emissions at each timestep based on output and carbon intensity, defined as the fossil fuel share in energy production. The CO2 emissions from various regions are input into the FAIR climate model, which aggregates them to calculate global mean surface temperature, expressed in °C above pre-industrial levels. This global temperature is then downscaled to regional temperature rise using a data-driven statistical downscaler. The downscaled regional temperature is used by the Damage Function to compute the fraction of economic output damaged by temperature increases in every region. Additionally, the Emissions sub-model allows for CO2 mitigation, with associated costs calculated by the Abatement module, based on the emission control rate chosen by agents (or regions) at each timestep. Detailed descriptions and equations for each model component are available in Appendix B 22footnotemark: 2.
Figure 1: Overview of Justice. The main outputs of the model have been highlighted with red arrows.
3.2 Multi-Agent Multi-Objective Formulation
We model Justice as a MOMAMDP with team reward as specified below.
Agents
We model 12 agents, each representing a macro-region that groups countries following RICE region specification (Nordhaus and Yang, 1996). A map of these macro regions with regional mapping table is provided in Appendix A 22footnotemark: 2. While Justice can generate data for 57 entities, modelling a MOMARL problem with 57 regions is computationally expensive and complicates analysis.
Actions
Each agent’s action is a two-dimensional discrete vector, where the actions of agent $n$ are:
  • Emission Control Rate (ECR) $\in\{0,0.1,0.2,\dots,1.0\}$,
  • Savings Rate (SR) $\in\{0,0.1,0.2,\dots,1.0\}$.
ECR represents the percentage of emissions removed from the atmosphere relative to the baseline SSP emissions. It is used in the Emissions module to assess emissions reduction, with the cost of this reduction being calculated in the Abatement module. SR determines the level of investment, indicating the amount of capital saved and reinvested into the economy, which in turn influences the economic output of each agent (a higher savings rate leads to higher economic output). Each agent undertakes these actions annually.
Observations
The agents can observe their local outcomes, the global temperature, and the mitigation rates set by all the agents at the previous timestep. For $n$-th agent at timestep $t$, the observation and their ranges are as follows:
${\Omega}_{t,n}={Y}^{GROSS}_{t,n}\cdot{\omega}_{t,n},\quad{\Omega}_{t,n}\in[0,\infty),$ (1)
where ${\Omega}_{t,n}$ is economic damage by the climate change in trillion USD adjusted to 2005 Purchasing Power Parity (PPP), ${Y}^{GROSS}_{t,n}$ is the gross economic output and ${\omega}_{t,n}$ is the damage fraction for agent $n$ at timestep $t$ and is calculated in the Damage module of Justice.
${\Lambda}_{t,n}$ $={\zeta}_{t,n}\cdot{\varepsilon}^{*}_{t,n}\cdot\left(\frac{{AC1}_{t,n}}{2}{ECR}_{t,n}^{2}+\frac{{AC2}_{t,n}}{5}{ECR}_{t,n}^{5}\right),$
${\Lambda}_{t,n}$ $\in[0,\infty),$ (2)
where ${\Lambda}_{t,n}$ is the abatement cost (cost for the regional economy to mitigate CO2), in trillion USD adjusted to 2005 PPP, ${\zeta}_{t,n}$ is the statistical correction factor, ${AC1}_{t,n}$ and ${AC2}_{t,n}$ are the region-specific abatement coefficients of agent $n$ at timestep $t$ calculated in the Abatement module of Justice.
${Y}^{NET}_{t,n}={Y}^{GROSS}_{t,n}-{\Omega}_{t,n}-{\Lambda}_{t,n},\quad{Y}^{NET}_{t,n}\in[0,\infty),$ (3)
where ${Y}^{NET}_{t,n}$ is the net economic output of agent $n$ at timestep $t$ in trillion USD adjusted to 2005 PPP and is calculated in the Economy module of Justice.
${\varepsilon}_{t,n}$ $={CI^{*}}_{t,n}\cdot{Y}^{GROSS}_{t,n}\cdot(1-{ECR}_{t,n})+{AFOLU}^{*}_{t,n},$
${\varepsilon}_{t,n}$ $\in[0,\infty),$ (4)
where ${\varepsilon}_{t,n}$ represents total emissions (Annual CO2 emissions in Gigatonne, GtCO2 per year), ${CI^{*}}_{t,n}$ is the carbon intensity (exogenous SSP data), ${AFOLU}^{*}_{t,n}$ is the Agriculture, Forestry, and Other Land Uses emissions (exogenous SSP data) of agent $n$ at timestep $t$ and is calculated in the Emissions module of Justice.
${GMT}_{t}=\textit{FAIR}\left(\sum_{n\in N}{\varepsilon}_{{t-1},n}\right),\quad{GMT}_{t}\in[0,\infty),$ (5)
where ${GMT}_{t}$ is the rise in the global average surface temperature since the pre-industrial era at timestep $t$ (in degrees Celsius) outputted by the FAIR model in the climate module.
$\begin{split}{RMT}_{t,n}&={TC1}_{t,n}+{TC2}_{t,n}\cdot{GMT}_{t},\\ {RMT}_{t,n}&\in[0,\infty),\end{split}$ (6)
where ${RMT}_{t,n}$ is the rise in the regional average surface temperature since the pre-industrial era (in degrees Celsius), ${TC1}_{t,n}$ and ${TC2}_{t,n}$ are the region-specific downscaler coefficients for agent $n$ at timestep $t$ and is calculated in the Downscaler of the Climate module.
$\begin{split}\mathbf{{VECR}_{t,n}}&=\{\text{ECR}_{t-1,1},\text{ECR}_{t-1,2},\dots,\text{ECR}_{t-1,N}\},\\ \mathbf{{VECR}_{t,n}}&\in\{0,0.1,0.2,\dots,1.0\}^{N},\end{split}$ (7)
where $\mathbf{{VECR}_{t,n}}$ is a vector of emission control rates adopted by the agents at the previous timestep that agent $n$ observes at timestep $t$.
Rewards
The rewards are modelled as team rewards, which assume collaboration among agents, where all agents receive the same return vector for executing the policy, reflecting their shared goal of improving global outcomes. This setup can be easily adjusted to individual rewards if needed. In our approach, each agent receives a two-dimensional vector of continuous rewards based on the following metrics:
  • Inverse Global Temperature (IGT):
    ${IGT}_{t,n}=1/{GMT}_{t},\quad{IGT}_{t,n}\in[0,1].$ (8)
  • Global Economic Output (GEO):
    ${GEO}_{t,n}=\sum_{n=1}^{N}{Y}^{NET}_{t,n},\quad{GEO}_{t,n}\in[0,\infty).$ (9)
In the RL setup, agents maximize their reward. Thus, we use the inverse of the Global Mean Temperature to reflect the goal of minimizing temperature rise. Similarly, we use the absolute value of the net economic output to reflect our aim of maximizing economic performance.
Starting State
The Justice simulation is initialized in 2015, using the SSP-2 data. SSP-2 is commonly used in IAM literature and it is the continuation of current emission trends into the future.
Episode Termination
Each episode consists of 285 timesteps, with each timestep representing one year, starting from 2015 and continuing until 2300. This extended timeframe accounts for the lag effects of the climate response, allowing the evaluation of damages not only for the current period but also for the centuries that follow.
4 Experimental Setup
We train our agents using Multi-Objective Multi-gent Proximal Policy Optimization (MOMAPPO)(Felten et al., 2024b), an extension of Multi-Agent PPO (MAPPO) (Yu et al., 2022) designed for multi-objective settings. MOMAPPO decomposes a multi-objective problem into multiple single-objective problems using a weighted-sum scalarization function for simplicity. Rewards are normalized to ensure consistency across objectives, and 100 weight vectors are uniformly sampled to generate diverse solutions. For each weight, MOMAPPO trains a multi-agent policy using MAPPO, evaluating its performance and adding non-dominated policies to the solution set. MOMAPPO is trained for 1 million global steps, with evaluations every (approximately) 20,000 steps for 10 random seeds. Appendix C 22footnotemark: 2 includes additional details.
4.1 Performance Indicators
Evaluating and comparing solution sets in MOMARL is more complex than in single-agent, single-objective RL due to the lack of inherent ordering of solution sets and intertwined performance across objectives. This added complexity leads to varied evaluation methods in MOMARL. To study convergence, we use two most commonly used approaches from both MORL and MARL domains for their suitability in MOMARL settings (Hayes et al., 2022) and to study the equality of distribution between agents we use GINI index:
Hypervolume ($\uparrow$)
(Zitzler et al., 2003) represents the region or (hyper-)volume between the points in the solution set and a reference point. The reference point indicates the lower bound for each objective. A solution set can be assessed by comparing its hypervolume with that of competing algorithms or the true Pareto front, if known.
Expected utility ($\uparrow$)
When the decision-maker’s utility function $u$ is linear, the expected utility (EU) metric (Vamplew et al., 2024) can be used to represent the expected utility over a distribution of reward weights $W$.
GINI Index ($\downarrow$)
The GINI index is a measure of inequality that quantifies disparities among agents based on a specific metric, with 0 indicating perfect equality and 1 indicating maximal inequality; we employ Concept-1 GINI by Milanovic (2011) to assess international inequality, where each agent represents a region and the index reflects whether their metrics are converging.
4.2 RICE50+ Comparison
We compare our results with the default RICE50+ model, as outlined in Section 2.1. This model consists of 57 regions and integrates optimization directly within the global climate simulation. The optimization follows a single-objective approach, employing a social welfare function as the global objective and using a single representative agent to optimize welfare, rather than a multi-agent multi-objective framework. The results presented in the next section are obtained using standard IAM methods, specifically non-linear programming within the General Algebraic Modeling Language (GAMS) for policy optimization. Unlike our approach, RICE50+ does not perform inter-temporal optimization at each timestep. Instead, it optimizes over the entire time horizon by using a social welfare function as a proxy for consumption per capita. Since consumption per capita is directly derived from net economic output (Equation 16a in Appendix B 22footnotemark: 2), we extract the net economic output from RICE50+ to compare with our objectives.
5 Results
The objectives of our experiments are to (1) verify the convergence of Justice MOMARL, (2) compare the solutions between Justice and the RICE50+ model, and (3) analyze key solutions from the Pareto set. We present results in line with these objectives.
5.1 Convergence
Figure 2 shows that our agents converge and demonstrate consistent training over time, with both performance metrics growing and eventually stabilizing. Note that both the hypervolume and Expected Utility metrics appear large due to their exponential scaling with the number of objectives and the achievable value ranges, particularly influenced by the high values of the GEO objective.
Figure 2: Mean hypervolume and expected utility over training steps (shaded area represents standard deviation).
5.2 Solutions
Figure 3 shows the Pareto set of solutions produced by Justice and the single RICE50+ solution (red star). We transform the objective values for simplicity: Total Global Economic Output represents the GEO objective, and Global Average Annual Temperature corresponds to the mean of the inverse of IGT (inverted to retrieve the temperature).
Figure 3: Pareto set of policies obtained by Justice (across 10 random seeds) with the RICE50+ polict for comparison. Arrows indicate the direction of preference for the objectives.
Justice produces 22 solutions (policies), shown as circles in Figure 3. Among these, three are highlighted: Climate Policy (purple), Economic Policy (green), and Compromise Policy (yellow). Although the purple policy is not the most extreme in terms of climate performance, it achieves substantial economic gains with only a slight increase in temperature compared to the extreme climate solution. Therefore, it is chosen as the Climate Policy. This example illustrates how the multi-objective approach supports decision-making by presenting a range of trade-offs.
The RICE50+ solution lies close to the Justice Pareto front, roughly in the middle, indicating a balance between economy and temperature rise. However, the Justice Compromise Policy dominates the RICE50+ solution, albeit by a small margin. Thus, Justice not only offers a range of solutions but also a slightly better solution than RICE50+.
5.3 Comparison of Key Solutions
We perform a comprehensive analysis of the highlighted policies in Figure 3—the three Justice policies and the RICE50+ policy. To do so, we plot the key IAM outcomes from 2015 to 2300 for these policies in Figure 4.
Figure 4: Performance of selected Justice policies over time (years on x-axis) for three important indicators: Global Net Economic Output, Temperature, and Cumulative Abated Emissions. The indicators show mean values and standard deviation (over 10 seeds) for selected Justice and RICE50+ policies over time.
The economic output
trajectories are shown in left panel of Figure 4. As expected, the best economic policy achieves the highest output. The best climate policy starts with the lowest output but eventually aligns with RICE50+ levels, as its rapid and deep mitigation efforts stabilize the climate and reduce damages, stimulating economic growth after 2150. The compromise policy presents an interesting trade-off; although it sacrifices slightly more output in the near term compared to RICE50+, it leads to significantly higher long-term growth, surpassing RICE50+ by hundreds of trillions of dollars, by reducing climate damages through immediate mitigation.
The temperature
trajectories are shown in the center panel of Figure 4. As expected, the best economic policy results in the highest temperature rise, while the best climate policy yields the lowest. Under the compromise policy, temperatures exceed 2°C by the end of the 2100s, reaching about 3.5°C by 2300. The RICE50+ temperature projections are significantly lower than those of the Justice compromise policy over the long term. This discrepancy highlights a common criticism of cost-benefit IAMs: their use of simple climate models inadequately captures the complexities and uncertainties of climate sensitivity, leading to smooth projections that miss abrupt, high-impact events (Mastrandrea and Schneider, 2001; Stanton et al., 2009; Pindyck, 2017). Such models tend to underestimate potential damages and future temperature rises, leading to policy inertia and delayed action (Füssel, 2006; Stern, 2022). We address these issues by using the FAIR model, which captures nonlinear dynamics and feedback mechanisms, effectively handling uncertainty in climate sensitivity (see Appendix B 22footnotemark: 2). FAIR produces temperature scenarios that align with more complex Earth System Models featured in the IPCC (O’Neill et al., 2016).
The abated emissions
trajectories are shown in the right panel of Figure 4. All three Justice policies favour rapid near-term mitigation, unlike RICE50+, which delays mitigation and peaks only around 2150. This finding aligns with the criticisms of traditional CB-IAMs for undervaluing early mitigation efforts. Among Justice policies, the best climate policy (purple trajectory) achieves the highest cumulative abated emissions. Both the best economic and compromise policies show similar mitigation levels, though the compromise policy emphasizes earlier action. This finding also aligns with the IPCC’s recommendations for rapid near-term mitigation to limit global warming and promote sustainable growth with minimal climate damage (Shukla et al., 2022).
5.4 Equity Analysis
Figure 5 shows the GINI index for emissions among 12 macro-regional agents, illustrating the distribution of mitigation and the future emissions budget. The emissions metric is used because it captures the agent’s primary action of emissions control. A basic setup in JUSTICE is employed without explicit equity objectives (such as Utilitarianism in RICE50+) to examine how disaggregating objectives and including multiple agents affect equity outcomes. For consistency, RICE50+ results from 57 regions have been aggregated to match our 12-region model.
Figure 5: GINI Index of Emissions Over Time for 12 Regions (years on x-axis).
As seen in Figure 5, the best climate and economic policies yield higher emissions inequality than the compromise policy. In the climate policy, inequality tends to be higher on average compared to other policies, as developing countries are potentially required to limit emissions despite their growth needs. In contrast, the GINI index for the economic policy is higher than the compromise because high-emitting developed countries are likely to increase emissions to drive economic growth. The compromise policy has the lowest average GINI index and a more equitable emissions distribution compared to RICE50+, which, despite a similar starting point, sees inequality rise sharply to around 0.5 by 2100. This setup demonstrates that the objectives that agents optimize and the chosen Pareto-optimal policy shape the allocation of mitigation burdens and emissions budgets across agents, setting the MOMARL approach apart from single-objective models. Fair policy design requires distinct equity objectives that convert outputs from Justice, such as regional damage and abatement costs, into an equity score while further disaggregating regions to better represent smaller nations. Nonetheless, the MOMARL approach is a promising step towards equitable policy design.
6 Discussion
Our results demonstrate how the optimization setup in IAMs can significantly affect the resulting climate policy recommendations. We compare Justice with the traditional single-objective optimization of RICE50+, a successor of widely used RICE IAM (Nordhaus and Yang, 1996) for studying optimal mitigation pathways. Our findings underscore the advantages of multi-objective optimization, which offers a range of solutions compared to single-objective optimization. We do not make any policy recommendations in this study; rather, we illustrate how policy outcomes are sensitive to the framing of the optimization problem and highlight how equity considerations vary with different solutions and objectives. This flexibility is crucial when agents have unequal economic capacities and face disproportionate climate impacts in an uncertain future.
Justice can have real-world impact in several ways.
Decision Support
The multi-objective multi-agent nature makes Justice a powerful decision-support tool for climate (e.g., COP) negotiations. In this use case, each Justice agent can represent a country or a coalition with associated preferences, simulating various policy options and implications. This enables stakeholders to propose, reassess, and refine actions, promoting transparent and robust negotiations.
Scientific Discourse
Our model can enrich scientific (e.g., IPCC) reports by providing complementary insights that capture dimensions often overlooked by conventional models. By exploring additional strategies through multi-objective multi-agent reinforcement learning, our approach broadens the scope of potential solutions and offers innovative pathways for regions to understand the consequences of their actions and implement more equitable climate policies.
Broader Engagement
While IAMs are crucial for policy recommendations, their technical nature and costly software licenses (e.g., models written in GAMS) often limit accessibility. By offering our framework as open source, we empower stakeholders to explore different scenarios and policy impacts. A more user-friendly interface could enhance this further. Further, our approach bridges research (e.g., IAM and AI) communities, demonstrating the benefits of building on models and algorithms developed by each other.
6.1 Limitations and Directions
The following are key limitations of our work. Addressing these limitations opens interesting avenues for future work.
Reward and Utility
The MOMARL algorithm we employ makes two key simplifying assumptions. First, a team-based reward structure, where all agents receive the same global reward. This is a simplification since, in practice, countries are likely to prioritize their individual interests. Second, it requires a linear utility function for the multi-objective aspect, which simplifies real-world scenarios. To our knowledge, this is the only MOMARL algorithm suitable for our setup of continuous state and reward spaces, and a discrete action space (see Table 2 in Felten et al. (2024b)). Relaxing these assumptions, i.e., exploring individual reward structures and other utility functions, is an important direction for future work.
Scalability
We perform analyses considering 12 macro-regions of the world. Although this is a substantial improvement over the single-agent assumption of traditional IAMs, a more fine-grained resolution may be necessary to derive insights on equity. This requires a substantial speedup of current MOMARL algorithms. One direction in this regard is to experiment with alternative weight-sampling techniques, such as adaptive weight sampling (Felten et al., 2024a).
Communication
Our current model assumes cooperation among regions, whereas real climate negotiations involve complex bargaining and conflicting agendas. Exploring different communication strategies among agents, and modeling trade and capital transfers, could improve the model’s ability to reflect real-world dynamics. Further, training the agents on different climate ensembles of FAIR, which represent varying plausible climate sensitivities, would allow the agents to take actions that are robust under climate uncertainty.
Policy Implementation
We acknowledge that the policies generated by Justice may be unrealistic, as both RICE50+ and Justice assume mitigation begins at the start of the model run, and the drastic mitigation suggested by Justice is infeasible due to geopolitical dynamics, technological inertia, and policy implementation challenges. However, the utility of Justice is in providing a range of alternatives, which can serve as a scientific basis for stakeholder deliberations.
User Studies
Finally, assessing which aspects of this framework are most valuable to stakeholders in COP negotiations is crucial. Conducting user studies with decision-makers can offer insights into their perceptions of the results and identify which elements most influence policymaking.
6.2 Conclusions
We introduce the Justice framework, the first to integrate IAM with MOMARL. Our approach stands apart from the models typically featured in climate reports—models upon which policymakers rely to inform decisions and negotiate actions. Unlike traditional models, which often oversimplify climate change into problems with a single agent and objective, our model embraces the complexity of the real world by acknowledging its multi-agent, multi-objective nature. Our experiments show that Justice produces flexible policies and allows detailed exploration of equity compared to RICE50+. Additionally, our model is open-source, implemented in Python, and designed to be accessible for exploration and experimentation of RL algorithms for an important real-world problem.
Appendix A Regions
Justice allows for 57 agents, which translate to 57 countries or regions. For computational efficiency we clustered the countries into 12 regions (Figure 6), as it was done in the regional RICE model (Table LABEL:tab:justice_macro_regions_grouped shows the detailed clustering).
Figure 6: Regions used in the MOMARL formulation
Table 1: Justice Regional Configuration: List of Macro Regions, Justice Regions, and Corresponding Countries
12 Macro Regions Justice Regions Countries
Latin America Argentina Argentina
Brazil Brazil
Chile Chile
Mexico Mexico
Rest Central America Anguilla, Antigua and Barbuda, Aruba, Bahamas, Barbados, Belize, Bermuda, Bonaire, Sint Eustatius and Saba, British Virgin Islands, Cayman Islands, Costa Rica, Cuba, Curaçao, Dominica, Dominican Republic, El Salvador, Grenada, Guatemala, Guadeloupe, Haiti, Honduras, Jamaica, Martinique, Montserrat, Nicaragua, Panama, Puerto Rico, Saint Barthelemy, Saint Kitts and Nevis, Saint Lucia, Saint Martin (French part), Saint Pierre and Miquelon, Saint Vincent and the Grenadines, Sint Maarten (Dutch part), South Georgia and the South Sandwich Islands, Trinidad and Tobago, Turks and Caicos Islands, United States Virgin Islands
Rest South America Bolivia, Colombia, Ecuador, Falkland Islands (Malvinas), French Guiana, Guyana, Paraguay, Peru, Suriname, Uruguay, Venezuela
Other High Income Australia Australia
Canada Canada
South Korea South Korea
Europe Austria Austria
Belgium Belgium
Baltic States Estonia, Latvia, Lithuania
Denmark Denmark
Spain Spain
Finland Finland
France France
United Kingdom United Kingdom
Greece Greece
Hungary Hungary
Ireland Ireland
Italy Italy
Netherlands Netherlands
Norway Norway
Poland Poland
Portugal Portugal
Czechia Czech Republic
Germany Germany
Slovakia Slovakia
Slovenia Slovenia
Switzerland Switzerland
Sweden Sweden
Turkey Turkey
Eurasia Bulgaria Bulgaria
Croatia Croatia
Other European Countries Åland Islands, Albania, Andorra, Bosnia and Herzegovina, Cyprus, Faroe Islands, Gibraltar, Greenland, Guernsey, Iceland, Isle of Man, Jersey, Kosovo, Liechtenstein, Luxembourg, Malta, Monaco, Montenegro, North Macedonia, San Marino, Serbia, Svalbard and Jan Mayen, Vatican City
Former Soviet Union Armenia, Azerbaijan, Belarus, Georgia, Kazakhstan, Kyrgyzstan, Moldova, Tajikistan, Turkmenistan, Uzbekistan
Romania Romania
Ukraine Ukraine
China China China
Japan Japan Japan
Russia Russia Russia
Middle East Egypt Egypt
Gulf Countries Bahrain, Iran, Iraq, Kuwait, Oman, Qatar, Saudi Arabia, United Arab Emirates, Yemen
Middle East Israel, Jordan, Lebanon, Palestine, Syria
Western Sahara, Tunisia, Morocco Morocco, Tunisia, Western Sahara
Algeria, Libya Algeria, Libya
Africa Sub-Saharan Africa Angola, Benin, Botswana, British Indian Ocean Territory, Bouvet Island, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Congo, Democratic Republic of the Congo, Côte d’Ivoire, Djibouti, Equatorial Guinea, Eritrea, Eswatini, Ethiopia, French Southern and Antarctic Lands, Gabon, Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Madagascar, Malawi, Mali, Mauritania, Mauritius, Mayotte, Mozambique, Namibia, Niger, Nigeria, Réunion, Rwanda, Saint Helena, São Tomé and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Sudan, Sudan, Tanzania, Togo, Uganda, Zambia, Zimbabwe
South Africa South Africa
India India India
Other Asia Indonesia Indonesia
Malaysia Malaysia
Other Southeast Asia Brunei Darussalam, Cambodia, Cocos (Keeling) Islands, Hong Kong, Lao People’s Democratic Republic, Macao, Mongolia, Myanmar, North Korea, Philippines, Singapore, Taiwan
Rest South Asia Afghanistan, Bangladesh, Bhutan, Maldives, Nepal, Pakistan, Sri Lanka
Thailand Thailand
Vietnam Vietnam
Rest Pacific Islands American Samoa, Christmas Island, Cook Islands, Fiji, French Polynesia, Guam, Heard Island and McDonald Islands, Kiribati, Marshall Islands, Micronesia, Nauru, New Caledonia, New Zealand, Niue, Norfolk Island, Northern Mariana Islands, Palau, Papua New Guinea, Pitcairn, Samoa, Solomon Islands, Tokelau, Tonga, Tuvalu, United States Minor Outlying Islands, Vanuatu, Wallis and Futuna
United States United States of America United States
Appendix B Additional Information on Justice
B.1 Economy Module
The Economy module employs the neoclassical Cobb-Douglas production function, as outlined in the established DICE and RICE integrated assessment models Nordhaus (2018) (see Equation 10). Here, the notation ${Y}^{GROSS}_{t,n}$ represents the gross economic output, ${TFP}^{*}_{t,n}$ represents exogenous Total Factor Productivity, and ${L}^{*}_{t,n}$ represents exogenous labor projections from the SSP dataset. The capital, denoted as ${K}_{t,n}$, is raised to the power of $\alpha$, and labor to the power of $1-\alpha$, where $\alpha$ signifies the capital elasticity in the production function. Capital dynamics at each time step are computed by considering the capital from the prior time step, adjusting for annual depreciation rate $\delta$, and incorporating the current time step’s investment, ${I}_{t,n}$, as detailed in Equation 10a.
${Y}^{GROSS}_{t,n}={{TFP}^{*}_{t,n}}\cdot{{K}}_{t,n}^{\alpha}\cdot{{L}^{*}_{t,n}}^{1-\alpha}$ (10)
${K}_{t,n}=(1-\delta)\cdot{K}_{{t-1},n}+{I}_{t,n}$ (10a)
B.2 Emissions Module
The economy model converts economic output into carbon dioxide emissions, ${\varepsilon}_{t,n}$, through the use of exogenous carbon intensity, ${CI^{*}}_{t,n}$, which reflects the proportion of fossil fuels in energy production (see Equation 11). Emissions from Agriculture, Forestry, and Other Land Uses (AFOLU) are also treated as exogenous, with values for both Carbon Intensity and AFOLU emissions, ${AFOLU}^{*}_{t,n}$, sourced from the SSP dataset. The emission control rate, ${ECR}_{t,n}$, serves as the model’s policy lever, determining the rate of mitigation and emissions for each time step.
$\begin{split}{\varepsilon}_{t,n}&={CI^{*}}_{t,n}\cdot{Y}^{GROSS}_{t,n}\cdot(1-{ECR}_{t,n})+{AFOLU}^{*}_{t,n}\end{split}$ (11)
B.3 Abatement Cost Module
Regions determine their abatement costs by using the emission control rate, ${ECR}_{t,n}$, which varies from 0 (no mitigation) to 1 (full mitigation), to mitigate a portion of baseline emissions. These costs are calculated using marginal abatement cost (MAC) curves, which are derived from the Enerdata dataset providing industrial carbon dioxide abatement levels at various carbon prices from 2025 to 2040. Beyond 2040, carbon price projections from the SSP database are used to create additional MAC curves. These MAC curves help calibrate region-specific coefficients, ${AC1}_{t,n}$ and ${AC2}_{t,n}$, alongside a correction factor, ${\zeta}_{t,n}$. Ultimately, the abatement cost correlates with the ${ECR}_{t,n}$ policy lever to compute the mitigation cost per time step, region, and state of the world, as outlined in Equation 12.
${\Lambda}_{t,n}={\zeta}_{t,n}\cdot{\varepsilon}^{*}_{t,n}\cdot\left(\frac{{AC1}_{t,n}}{2}{ECR}_{t,n}^{2}+\frac{{AC2}_{t,n}}{5}{ECR}_{t,n}^{5}\right)$ (12)
B.4 Climate Module
The Justice framework incorporates the established Finite-amplitude Impulse Response (FAIR) climate module, a simplified model designed to replicates the behavior of more complex Earth System Models featured in the Coupled Model Intercomparison Project Phase 6 (CMIP6) (O’Neill et al., 2016). FAIR employs a set of six equations to efficiently simulate the global mean climate response to 52 greenhouse gases (GHGs) over the period from 1750 to 2100 (Leach et al., 2021). Total CO2 emissions, encompassing those from fossil fuels, industrial processes modeled by Justice IAM, and emissions from agriculture, forestry, and other land use (AFOLU), are used to estimate GHG concentrations. Other GHG emissions, as well as data on volcanic activity and solar forcing, are treated as exogenous inputs sourced from the Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP) datasets via the Reduced Complexity Model Intercomparison Project (RCMIP).
The FAIR model’s structure includes five equations that compose the standard impulse model utilized for GHG metrics in the IPCC’s Fifth Assessment Report, plus an additional equation that integrates state-dependent feedback, addressing nonlinearities in the carbon cycle due to atmospheric decay interactions within carbon and methane cycles. This model functions by converting GHG emissions into concentrations, which are subsequently used to calculate effective radiative forcing, culminating in the computation of temperature response by the climate system. The entire process is primarily divided into three components: the gas cycle, radiative forcing, and climate response (See Figure 7).
FAIR’s computational efficiency enables robust exploration of climate uncertainty through a constrained and calibrated ensemble, where uncertainty is quantified in terms of Equilibrium Climate Sensitivity (ECS) and Transient Climate Response (TCR). The initial ensemble consisted of 1.5 million members, capturing uncertainties within the carbon cycle, effective radiative forcing, and thermal responses, aligning with IPCC’s estimated ECS and TCR ranges and historical climate data, covering ocean heat uptake, global mean surface temperature, aerosol radiative forcing, carbon dioxide levels, and future warming forecasts from the SSP2-RCP4.5 scenario. This full ensemble was narrowed down to a constrained sample of 1,001 members based on each member’s alignment with the likelihood of present-day warming. Each member of this constrained ensemble successfully reproduces historical warming trends, representing distinct potential “states of the world”, (Smith et al., 2023). FAIR computes the global mean temperature increase, ${GMT}_{t}$, at each time step $t$ by integrating regional emissions from the previous time step $t-1$, according to Equation 13.
${GMT}_{t}=\textit{FAIR}\left(\sum_{n\in N}{\varepsilon}_{{t-1},n}\right),\quad\forall t\in T,\ \forall n\in N$ (13)
Figure 7: Overview of FAIR Climate Model
B.5 Regional Temperature Downscaler
Justice employs the climate downscaler from RICE50+ to determine the regional temperature response across 57 regions. This downscaler utilizes a statistical technique to evaluate temperature data from 244 countries, leveraging information from the Coupled Model Intercomparison Project to link global and local temperature changes via historical data. It integrates global temperature data from all Representative Concentration Pathways (RCPs), which have been implemented by various global climate models for future forecasts. Utilizing this dataset, a linear regression downscaler is calibrated with region-specific coefficients ${TC1}_{t,n}$ and ${TC2}_{t,n}$ to estimate the regional temperature in a given region $n_{j}$ at time step $t_{i}$, based on the global mean temperature rise ${GMT}_{t,n}$ as presented in Equation 14. Justice then distributes the global mean temperature rise to all regions in set $N$ to derive ${GMT}_{t,n}$, as outlined in Equation 14a.
${RMT}_{t,n}={TC1}_{t,n}+{TC2}_{t,n}\cdot{GMT}_{t,n}$ (14)
$\text{where}\quad gmt_{t,n}=gmt_{t},\quad\forall n$ (14a)
B.6 Climate Damage Module
Justice uses the RICE50+ model’s methodology to estimate the damage fraction, ${\omega}_{t,n}$, which quantifies the portion of economic output lost to climate damages (see Equation 15). This fraction is derived from the impact factor, ${\psi}_{t,n}$, calculated using the impact from the previous time step, $t-1$, and the damage function specification, ${\sigma}_{t,n}$, through a simplified impact formulation (Equation 15a). The damage specification follows Kalkuhl and Wenz’s approach (Kalkuhl and Wenz, 2020), analyzing data from over 1,500 regions in 77 countries. This method utilizes subnational GDP data (Gross Regional Product, GRP) to empirically estimate historical climate impacts over various timeframes and regions. It captures linear impacts on production based on lagged temperature, ${RMT}_{t-1}$, and accounts for weather changes by including an interaction term: the product of the annual regional temperature change, $\Delta{RMT}_{t}$, and the lagged temperature, ${RMT}_{t-1}$. The annual temperature-dependent impact on productivity is calculated using coefficients ${DC1}_{t}$ and ${DC2}_{t}$, which are applied to both the lagged temperature and the interaction term (Equation 15b).
${\omega}_{t,n}=1-\frac{1}{1+{\psi}_{t,n}}$ (15)
${\psi}_{t,n}=\frac{1+{\psi}_{t-1,n}}{1+{\sigma}_{t-1,n}}-1$ (15a)
${\sigma}_{t,n}={DC1}_{t,n}\cdot\Delta{RMT}_{t,n}+{DC2}_{t,n}\cdot{RMT}_{t-1,n}\cdot\Delta{RMT}_{t,n}$ (15b)
$\text{where}\quad dc1_{t,n}=dc1\quad\forall t,n$ (15c)
$\text{and}\quad dc2_{t,n}=dc2\quad\forall t,n$ (15d)
$\Delta{RMT}_{t,n}={RMT}_{t,n}-{RMT}_{t-1,n}\quad$ (15e)
B.7 Consumption Calculation
Consumption, represented as ${C}_{t,n}$, for a region at a given time step and state, is calculated based on its net economic output, ${Y}^{NET}_{t,n}$, and investment, ${I}_{t,n}$, as described in Equation 16. To determine consumption per capita, the consumption hypermatrix is divided by the population or labor ${L}^{*}_{t,n}$ from the SSP database (Equation 16a). Net economic output is derived by deducting climate impact costs, ${\Omega}_{t,n}$, and abatement costs, ${\Lambda}_{TNS}$, from the gross economic output, which accounts for the costs of mitigating CO2 emissions (Equation 16b). The damage cost is computed by multiplying the gross economic output, ${Y}^{GROSS}_{t,n}$, by the damage fraction (Equation 16c). Investment ${I}_{t,n}$ signifies the portion of ${Y}^{GROSS}_{t,n}$ that is saved and reinvested to drive economic growth. This saving depends on the savings rate ${S}_{t,n}$, unique to each region per time step, which is multiplied by the net economic output to determine investment (Equation 16d). In this study, the savings rate is exogenous, sourced from regional historical data, and aligns with long-term projections from DICE-2016R2 (Nordhaus, 2018).
${C}_{t,n}={Y}^{NET}_{t,n}-{I}_{t,n}$ (16)
${CPC}_{t,n}=\frac{{C}_{t,n}}{{L}^{*}_{t,n}}$ (16a)
${Y}^{NET}_{t,n}={Y}^{GROSS}_{t,n}-{\Omega}_{t,n}-{\Lambda}_{t,n}$ (16b)
${\Omega}_{t,n}={Y}^{GROSS}_{t,n}\cdot{\omega}_{t,n}$ (16c)
${I}_{t,n}={S}_{t,n}\cdot{Y}^{NET}_{t,n}$ (16d)
Appendix C MOMAPPO Setup
We train our agents using MOMAPPO (Felten et al., 2024b), an extension of Multi-Agent PPO (MAPPO) (Yu et al., 2022) designed for multi-objective settings. MOMAPPO generates a solution set of multi-agent policies for cooperative problems by leveraging a decomposition approach (Felten et al., 2024a). This technique breaks the multi-objective problem into a collection of single-objective problems, which are then solved using a multi-agent reinforcement learning algorithm. A scalarisation function, parameterised by weight vectors, is employed to perform the decomposition and focus on different regions of the objective space. In our work, we use the weighted sum scalarisation function due to its simplicity. To ensure consistency across objectives, the environment rewards are first normalised to account for differences in scale. Weight vectors are generated by sampling 100 weights uniformly. For each sampled weight, MOMAPPO trains a multi-agent policy to optimise a specific trade-off using MAPPO. After training, the policy is evaluated on the original environment to estimate its value vector $\mathbf{v}^{\pi}$. If the policy is non-dominated, it is added to the solution set. The algorithm concludes by returning all non-dominated multi-agent policies, providing a diverse set of solutions for various trade-offs in the objective space.
We train MOMAPPO for 1 million global steps, evaluating the performance every (approximately) 20,000 steps. The policies are implemented as Multi-Layer Perceptron (MLP) networks consisting of two layers, each with 64 hidden dimensions, and using Rectified Linear Unit (ReLU) activation functions. The training and evaluation is conducted across ten random seeds.